OpenStudy (lasttccasey):

A buoy oscillates in simple harmonic motion y = A cos(ωt) as waves move past it. The buoy moves a total of 4.5 feet (vertically) from its low point to its high point. It returns to its high point every 2 seconds. (a) Write an equation describing the motion of the buoy if it is at its high point at t = 0. y= (b) Determine the velocity of the buoy as a function of t. v=

OpenStudy (lasttccasey):

No idea where to start with the equation..

OpenStudy (swissgirl):

well lets start with determing teh equation

OpenStudy (swissgirl):

So we know that from the max to the min it travles all of 4.5 feet. We know that the amplitude wich is represented by the A is 1/2 of the total amt so what is the amplitude?

OpenStudy (lasttccasey):

2.25 of course but the ω in the cos is throwing me off

OpenStudy (swissgirl):

Yes ok next lets decide the period which is that funny looking w

OpenStudy (swissgirl):

we know that the period is 2 seconds

OpenStudy (swissgirl):

since it reaches its maximum every 2 seconds

OpenStudy (lasttccasey):

omega i believe its called, y= 2.25cos(πt)?

OpenStudy (lasttccasey):

then take the derivative of that and I can find velocity right?

OpenStudy (swissgirl):

lets use k instead of that funny w cuz its neways i dummy variable so there is a formula 360/k=period our period in this case is 2 360/k=2

OpenStudy (swissgirl):

360/2=k k=180 so our funny w is equal to 180 did u follow?

OpenStudy (lasttccasey):

did you see my reply? lol

OpenStudy (swissgirl):

Well do u have the equation??? First u gotta determine the equation and then u can determine the velocity

OpenStudy (swissgirl):

If u want i can give u the equation if u dont want to understand how i got he equation i dont mind to be honest with you

OpenStudy (swissgirl):

y=2.25Cos(180t) Now u can find the derivative

OpenStudy (lasttccasey):

I have the equation.. it was in my reply but yes I do understand. 2.25cos(πt)

OpenStudy (lasttccasey):

we have to do everything in terms of π but thanks.

OpenStudy (swissgirl):

ohhhhhh even the the omega must???

OpenStudy (swissgirl):

that is weird ohh interesting sorry for wasting ur time i didnt realize that

OpenStudy (lasttccasey):

no its okay

OpenStudy (swissgirl):

U r ok with the rest right? hahah ik y u got annoyed at me. it happens to me too. Like i dont know like how much u knew and explained the topic as if u didnt get it at all. Ppl do it to me also.

OpenStudy (lasttccasey):

yea, it happens to me too. Hard to tell how much people know but this is for calc I so I better know this stuff lol. When I did it I dropped the π by accident but its correct now. Thanks.

OpenStudy (swissgirl):

hahahah i finished calc 3 and cant remember calc 1 lol